A rectangle with a width of (-5x + 4) feet
and a length of (-6x +7) feet, has an area of
100 square feet. What is the rectangle's
width and length?

Answer by stanbon(75887)

A rectangle with a width of (-5x + 4) feet
and a length of (-6x +7) feet, has an area of
100 square feet. What is the rectangle's
width and length?
------
Equation:
(-5x+4)(-6x+7) = 100
-----
30x^2 -35x-24x + 28 = 100
----
30x^2 - 57x - 72 = 0
-----
x = 2.767
-----
width = -5x+4 = a negative value
---
width cannot be negative.
Pls check your post.
Cheers,
Stan H.
----------